# A discrete Consensus-Based Global Optimization Method with Noisy Objective Function

### v1

@@ Abstract
Keywords: global optimization, consensus-based optimization, noisy-functions, subsampling.
@@ 1 Introduction
Global optimization problems are widespread in mathematical modeling of real problems and a large variety of computational approaches have been proposed for their solution. For an introduction on the topic, see e.g. [14, 16].
@@
where $\gamma\in(0,1]$, $\sigma>0$ are the drift and the diffusion parameter respectively, $\{\{B^{i}_{t}\}_{t\geq 0}|i=1,\dots,N\}$ are independent Brownian motions in $\mathbb{R}^{d}$ and for a given value of $\alpha>0$ the consensus point ${x}^{\alpha}_{k}$ is defined as
In [6] the authors perform the theoretical analysis of the method described in (2) by studying the evolution of its mean-field limit. First of all, they prove that, under suitable regularity assumptions on the objective function, the particles achieve consensus. More precisely, they prove that, asymptotically, the distribution of the mean-field limit tends to a distribution centered at one point. Finally, they provide assumptions on the initial distribution of the particles, that ensure that the consensus point is in a neighborhood of the global minimizer of the objective function, where the size of the neighborhood depends on the parameter $\alpha.$ In [11] the authors prove that the probability distribution of the mean-field limit of the particles converges to a distribution centered at the solution of the problem, in the Wasserstein distance, with minimal assumptions on the initial distribution of the particles.
@@
where $\{\eta_{k,s}\}$ are i.i.d. random variables with mean zero and variance $\sigma^{2}$.
@@ 3 Convergence Analysis
Moreover, $f$ is locally strictly convex at ${x}^{*}$. That is, $\nabla^{2}f({x}^{*})$ is positive definite.
In the following we define $f_{*}=f({x}^{*}).$
@@
where $E_{\omega({x})}$ denotes the expected value with respect to the distribution of the random variable $\omega({x})$ given ${x}$. If the oracle is unbiased (15) is equivalent to
@@
We assume that the variance of the function estimate satisfies a growth condition. This is weaker and more realistic than the bounded variance condition and was recently used in the analysis of stochastic optimization methods, see [4, 8].
We start by studying the asymptotic behavior of the generated sequences $\{x_{k}^{i}\}$, $i=1,\ldots,N$. Namely, we prove that with probability one, all the particles converge to the same point ${x}_{\infty}$, as $k$ tends to +.$\infty$.+ ∞ . The proof of convergence follows the path of [13]. We fist provide some auxiliary results.
@@ 3.1 Diameter and Ergodicity
    $\lim_{k\rightarrow+\infty}S_{k,s}=1-{\bf\vartheta}\enspace\text{almost surely}$
@@
$\lim_{k\rightarrow+\infty}S_{k,s}={\mathbb{E}}\left[\gamma-4\max_{i=1:N}|\eta^{i}_{l,s}|\right]\leq\gamma-8\xi\log(\sqrt{2}N)^{1/2}=1-{\bf\vartheta}\enspace\text{almost surely},$
@@
$\leq$ $\max_{s=1:d}e^{-kS_{k,s}}\operatorname{diam}({y}^{s}_{0}){\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\leq}\max_{s=1:d}e^{-kS_{k,s}}\max_{s=1:d}\max_{i,j=1:N}|x^{i}_{0,s}-x^{j}_{0,s}|$
@@ 3.2 Almost sure consensus
Let ${x}_{k}^{i}$, $i=1,\ldots N$ be the $N$ particles generated at iteration $k$ of (7), ${y}_{k}$ given in (12) and ${\bf\vartheta}$ given in (10). If $\gamma\in(0,1)$, $\xi>0$ and $N\in\mathbb{N}$ are such that ${\bf\vartheta}<1$, then
@@
By Lemma 3.2 we have ${\mathbb{E}}[\operatorname{diam}({y}^{s}_{k})]=\mathcal{O}\left({\bf\vartheta}^{k}\right)$, $\operatorname{diam}({y}^{s}_{k})=\mathcal{O}\left(e^{-k(1-{\bf\vartheta})}\right)$ and $\max_{i,j=1:N}\|{x}^{i}_{k}-{x}^{j}_{k}\|_{\infty}=\mathcal{O}\left(e^{-k(1-{\bf\vartheta})}\right)$ almost surely. Then, the thesis follows immediately since ${\bf\vartheta}<1$. ∎
@@
Part $iii)$ follows from the fact that $\nabla f$ is continuous and, by $ii)$, the sequence $\{\hat{x}^{\alpha}_{k}\}_{k=0}^{\infty}$ is bounded. with probability 1. ∎
@@ 3.3 Handling the noise
Given $\{x^{1}_{k,s},\dots,x^{N}_{k,s}\}\in\mathbb{R},\alpha>0$ we define the function $\Psi_{k,s}:\mathbb{R}^{N}\longrightarrow\mathbb{R}$
@@
which concludes the proof. ∎
@@
which yields the thesis with $M_{v}=N^{1/2}\left(t_{0}+t_{1}M_{f}^{2}\right)^{1/2}$. Let us now consider the case of additive Gaussian noise. Using (25), the tower rule, and Lemma A.1 in the Appendix we have
@@ 3.4 Almost sure convergence
for some $\theta^{i}_{k}\in(0,1).$ Taking the average over $i=1,\dots,N$ we get the thesis. ∎
@@
By Lemma 3.6 and 3.7, inequalities (19) and (20) we have for $s=1,\dots,d$:
@@
and part $iv)$ of Lemma 3.2 yields the thesis. ∎
@@
[12, Proposition 3.1] Let Assumptions A1, A3 and A4 hold, and let ${x}_{0}$ denote a random variable with law $\rho_{0}$ given in Assumption 4. Then
@@
$\left|\operatorname*{ess\,inf}f({x}_{\infty})-f_{*}\right|\leq\frac{d}{2}\frac{\log(\alpha)}{\alpha}+\mathcal{O}\left(\frac{1}{\alpha}\right),$
@@
Taking the expectation in (28), using Lemma 3.9 and 3.10 we have
@@
$\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[e^{-\alpha f({x}^{i}_{k+1})}\right]$ $\geq\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[e^{-\alpha f({x}^{i}_{0})}\right]-\alpha e^{-\alpha f_{*}}\left(\Gamma_{A}\sum_{l=0}^{k}\zeta_{l}^{2}+\Gamma_{B}\sum_{l=0}^{k}\zeta_{l}\right)$
@@
$\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[e^{-\alpha f({x}^{i}_{0})}\right]=\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[e^{-\alpha f({x}_{0})}\right]={\mathbb{E}}\left[e^{-\alpha f({x}^{i}_{0})}\right].$ (36)
@@
$=f_{*}+\frac{d\log(\alpha)}{2\alpha}+\mathcal{O}\left(\frac{1}{\alpha}\right)$
which gives the thesis. ∎
As noticed in [12], Theorem 3.12 provides an error estimate only for small values of $\alpha$, as when $\alpha$ increases, the condition (35) becomes very restrictive.
@@ 3.5 Convergence in expectation
$+2(\gamma^{2}+\xi^{2}+(\gamma^{2}+\xi^{2})^{1/2}){\mathbb{E}}[V_{k}]^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}$
@@
$Z_{k,1}=\sum_{i=1}^{N}\|(\Delta x)_{k}^{i}\|^{2}=\frac{1}{N}\sum_{i=1}^{N}\left\|(\gamma I_{d}+{\bf\Xi}_{k}^{i})(x^{i}_{k}-\hat{x}^{\alpha}_{k})\right\|^{2},$ (43)
@@
$Z_{k,1}$ $\leq\frac{1}{N}\sum_{i=1}^{N}\left\|(\gamma I_{d}+{\bf\Xi}_{k,s}^{i})(x^{i}_{k}-{x}^{*})\right\|^{2}$ (44)
$+\frac{2}{N}\sum_{i=1}^{N}\left\|(\gamma I_{d}+{\bf\Xi}_{k}^{i})(x^{i}_{k}-{x}^{*})\right\|\left\|(\gamma I_{d}+{\bf\Xi}_{k}^{i})(\hat{x}^{\alpha}_{k,}-{x}^{*})\right\|$
@@
$=(\gamma^{2}+\xi^{2}){\mathbb{E}}[\|x^{i}_{k}-{x}^{*}\|^{2}]^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}$
Taking the expected value on both sides of (44) we then have
@@
$\quad+\frac{2}{N}(\gamma^{2}+\xi^{2})\sum_{i=1}^{N}{\mathbb{E}}[\|x^{i}_{k}-{x}^{*}\|^{2}]^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}_{s}\|^{2}]^{1/2}$
@@
$=(\gamma^{2}+\xi^{2})({\mathbb{E}}[V_{k}]+2{\mathbb{E}}[V_{k}]^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}$
@@
$+2(\gamma^{2}+\xi^{2})^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}[\|x^{i}_{k}-x^{*}\|^{2}]^{1/2}.$
Then, noting that, given $\{a_{i}\}_{i=1}^{N}$ with $a_{i}\geq 0$ for any $i$, it holds
we obtain
$\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}[\|x^{i}_{k}-x^{*}\|^{2}]^{1/2}\leq\frac{1}{\sqrt{N}}(\sum_{i=1}^{N}{\mathbb{E}}[\|x^{i}_{k}-x^{*}\|^{2}])^{1/2}={\mathbb{E}}[V_{k}]^{1/2}$
and by (47)
${\mathbb{E}}[Z_{k,2}]\leq-2\gamma{\mathbb{E}}[V_{k}]+2(\gamma^{2}+\xi^{2})^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}{\mathbb{E}}[V_{k}]^{1/2}.$ (49)
From (42), (45) and (49) we get the thesis. ∎
@@
Let us now derive an upper bound on $\|{x}_{k}^{\alpha}-{x}^{*}\|$, where ${x}_{k}^{\alpha}$ has been defined in (3). Consider $r\in(0,R_{0}]$ and $q>0$ such that $q+f_{r}\leq f_{\infty}$, with
@@
We distinguish the case $|{\cal I}_{k,r}|>0$ and $|{\cal I}_{k,r}|=0$. We stress that in the first case, at iteration $k$ there exists at least one particle that belongs to $B_{r}$.
Lemma 3.14.
@@
    $\|{x}^{\alpha}_{k}-{x}^{*}\|\leq\frac{1}{\beta}(q+f_{r})^{\nu}+\frac{e^{-\alpha q}}{|{\cal I}_{k,r}|}\sum_{i=1}^{N}\|{x}^{i}_{k}-{x}^{*}\|$
@@
Proof.
$\|{x}^{\alpha}_{k}-{x}^{*}\|$ $=\left\|\frac{1}{\sum_{i=1}^{N}w_{i,k}^{\alpha}}\sum_{i=1}^{N}w^{\alpha}_{i,k}({x}_{k}^{i}-{x}^{*})\right\|$ (53)
$\leq\frac{1}{\sum_{i=1}^{N}w_{i,k}^{\alpha}}\sum_{i=1}^{N}w^{\alpha}_{i,k}\left\|{x}_{k}^{i}-{x}^{*}\right\|$
By assumption A6 and the definition of $\tilde{r}$ we have
@@
Moreover, ${\mathbb{E}}[V_{k}]\leq\mu^{k}{\mathbb{E}}[V_{0}]$ for $k=0,\dots,k_{*}.$
@@
$\|\hat{x}^{\alpha}_{\bar{k}}-{x}^{*}\|^{2}=$ $\|{x}^{\alpha}_{\bar{k}}-{x}^{*}\|^{2}+\|{x}^{\alpha}_{\bar{k}}-\hat{x}^{\alpha}_{\bar{k}}\|^{2}+2\|{x}^{\alpha}_{\bar{k}}-{x}^{*}\|\|{x}^{\alpha}_{\bar{k}}-\hat{x}^{\alpha}_{\bar{k}}\|$ (63)
$\leq$ $\frac{1}{4}C_{\bar{k}}^{2}+N^{2}e^{-2\alpha q}V_{\bar{k}}+Ne^{-\alpha q}V_{\bar{k}}^{1/2}C_{\bar{k}}$
$+\alpha^{2}\|{\cal E}_{\bar{k}}\|_{\infty}^{2}\sum_{s=1}^{d}\operatorname{diam}({y}^{s}_{0})^{2}\prod_{l=0}^{{\bar{k}}-1}(1-\gamma+4\max_{i=1:N}|\eta_{l,s}^{i}|)^{2}$
$+C_{\bar{k}}\alpha\|{\cal E}_{\bar{k}}\|_{\infty}\left(\sum_{s=1}^{d}\operatorname{diam}({y}^{s}_{0})^{2}\prod_{l=0}^{{\bar{k}}-1}(1-\gamma+4\max_{i=1:N}|\eta_{l,s}^{i}|)^{2}\right)^{1/2}$
$+2\alpha Ne^{-\alpha q}V_{\bar{k}}^{1/2}\|{\cal E}_{\bar{k}}\|_{\infty}\left(\sum_{s=1}^{d}\operatorname{diam}({y}^{s}_{0})^{2}\prod_{l=0}^{{\bar{k}}-1}(1-\gamma+4\max_{i=1:N}|\eta_{l,s}^{i}|)^{2}\right)^{1/2}.$
@@
which concludes the proof. ∎

### v2

@@ 1 Introduction
Global optimization problems are widespread in the mathematical modeling of real problems and a large variety of computational approaches have been proposed for their solution. For an introduction on the topic, see e.g. [15, 17, 21, 23].
@@
where $\gamma\in(0,1]$, $\sigma>0$ are the drift and the diffusion parameter respectively, $\{\{B^{i}_{t}\}_{t\geq 0}|i=1,\dots,N\}$ are independent Brownian motions in $\mathbb{R}^{d}$ and for a given value of $\alpha>0$ the consensus point ${x}^{\alpha}_{t}$ is defined as
${x}^{\alpha}_{t}=\frac{\sum_{i=1}^{N}{x}^{i}_{t}e^{-\alpha f({x}^{i}_{t})}}{\sum_{i=1}^{N}e^{-\alpha f({x}^{i}_{t})}}.$
In [6] the authors perform the theoretical analysis of the method described in (2) by studying the evolution of its mean-field limit. First of all, they prove that, under suitable regularity assumptions on the objective function, the particles achieve consensus. More precisely, they prove that, asymptotically, the distribution of the mean-field limit tends to a distribution centered at one point. Finally, they provide assumptions on the initial distribution of the particles, that ensure that the consensus point is in a neighborhood of the global minimizer of the objective function, where the size of the neighborhood depends on the parameter $\alpha.$ In [11] the authors prove that the probability distribution of the mean-field limit of the particles converges to a distribution centered at the solution of the problem, in the Wasserstein distance, with minimal assumptions on the initial distribution of the particles. Moreover, they provide a probabilistic convergence result in the finite-particles regime. Specifically, they prove that the squared norm of the difference between the average particle and the solution becomes smaller than any given threshold with a certain probability.
@@
where $\{\eta_{k,s}\}$ are i.i.d. random variables with mean zero and variance $\sigma^{2}$, and the consensus point ${x}^{\alpha}_{k}$ is defined as
@@ 3 Convergence Analysis
Moreover, $f$ is locally strictly convex at ${x}^{*}$. That is, $\nabla^{2}f({x}^{*})$ is positive definite. In the following we define $f_{*}=f({x}^{*}).$
@@
where $E_{\omega({x})}$ denotes the expected value with respect to the distribution of the random variable $\omega({x})$ given ${x}$.
Remark 3.1.
Assumption A4 requires the unknown solution $x^{*}$ to be in the support of the initial distribution $\rho_{0}$ of the particles. This can be achieved, in practice, taking a distribution $\rho_{0}$ with support over the whole space $\mathbb{R}^{n}$. This assumption is necessary for our theoretical analysis in the discrete-time finite-particles regime. In [12] the authors prove, in the continuous-time setting, that the convergence of CBO method to the global minimizer can be proved also when the initial distribution has no mass around $x^{*}.$
Remark 3.2.
Assumption A5 does not require the variance of the oracle to go to zero and also applies to the setting in which available function value estimates are occasionally biased. If the oracle is unbiased,(15) is equivalent to
@@
Then we assume that the variance of the function estimate satisfies a growth condition. This is weaker and more realistic than the bounded variance condition and it is used in the analysis of stochastic optimization methods, see e.g [4, 8]. Note that in particular Assumption A5 does not require the expectation of the squared error to vanish as the true value of the function approaches 0.
We start by studying the asymptotic behavior of the generated sequences $\{x_{k}^{i}\}$, $i=1,\ldots,N$. Namely, we prove that with probability one, all the particles converge to the same point ${x}_{\infty}$, as $k$ tends to +.$\infty$.+ ∞ . The proof of convergence follows the path of [14]. We first provide some auxiliary results.
@@ 3.1 Diameter and Ergodicity
    $\lim_{k\rightarrow+\infty}S_{k,s}\geq 1-{\bf\vartheta}\enspace\text{almost surely}$
@@
$\lim_{k\rightarrow+\infty}S_{k,s}={\mathbb{E}}\left[\gamma-4\max_{i=1:N}|\eta^{i}_{l,s}|\right]\geq\gamma-8\xi\log(\sqrt{2}N)^{1/2}=1-{\bf\vartheta}\enspace\text{almost surely},$
@@
$\leq$ $\max_{s=1:d}e^{-kS_{k,s}}\operatorname{diam}({y}^{s}_{0})\leq\max_{s=1:d}e^{-kS_{k,s}}\max_{s=1:d}\max_{i,j=1:N}|x^{i}_{0,s}-x^{j}_{0,s}|$
@@ 3.2 Almost sure consensus
Let $\{({x}^{i}_{k})_{i=1:N}\}_{k}$ be the sequence generated by (7),${\bf\vartheta}$ given in (10), and ${y}_{k}^{s}$, $s=1,\ldots,d$ given in (12). If $\gamma\in(0,1)$, $\xi>0$ and $N\in\mathbb{N}$ are such that ${\bf\vartheta}<1$, then
@@
By Lemma 3.4 we have ${\mathbb{E}}[\operatorname{diam}({y}^{s}_{k})]=\mathcal{O}\left({\bf\vartheta}^{k}\right)$, $\operatorname{diam}({y}^{s}_{k})=\mathcal{O}\left(e^{-k(1-{\bf\vartheta})}\right)$ and $\max_{i,j=1:N}\|{x}^{i}_{k}-{x}^{j}_{k}\|_{\infty}=\mathcal{O}\left(e^{-k(1-{\bf\vartheta})}\right)$ almost surely. Then, the thesis follows immediately since ${\bf\vartheta}<1$. ∎∎
@@
Part $iii)$ follows from the fact that $\nabla f$ is continuous and, by $ii)$, the sequence $\{\hat{x}^{\alpha}_{k}\}_{k=0}^{\infty}$ is bounded. with probability 1. ∎∎
@@ 3.3 Handling the noise
Given $x^{i}_{k,s}\in\mathbb{R}$, $i=1,\ldots,N$, $\alpha>0$ we define the function $\Psi_{k,s}:\mathbb{R}^{N}\longrightarrow\mathbb{R}$
@@
which concludes the proof. ∎∎
@@
which yields the thesis. Let us now consider the case of additive Gaussian noise. Using (25), the tower rule, and Lemma A.1 in the Appendix we have
@@ 3.4 Almost sure convergence
for some $\theta^{i}_{k}\in(0,1).$ Taking the average over $i=1,\dots,N$ we get the thesis. ∎∎
@@
By Lemma 3.8 and 3.9, inequalities (19) and (9) we have for $s=1,\dots,d$:
@@
and part $iv)$ of Lemma 3.4 yields the thesis. ∎∎
@@
[13, Proposition 3.1] Let Assumptions A1, A3 and A4 hold, and let ${x}_{0}$ denote a random variable with law $\rho_{0}$ given in Assumption 4. Then, for $\alpha$ sufficiently large,
@@
$\left|\operatorname*{ess\,inf}f({x}_{\infty})-f_{*}\right|<\frac{d}{2}\frac{\log(\alpha)}{\alpha}+G(\alpha)$ (36)
for some function $G(\alpha)=\mathcal{O}\left(\frac{1}{\alpha}\right)$.
@@
Taking the expectation in (28), using Lemma 3.11 and 3.12 we have
@@
$\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[e^{-\alpha f({x}^{i}_{k+1})}\right]$ $\geq\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[e^{-\alpha f({x}^{i}_{0})}\right]-\alpha e^{-\alpha f_{*}}\left(\Gamma_{A}\sum_{l=0}^{k}\zeta_{l}^{2}+\Gamma_{B}\sum_{l=0}^{k}\zeta_{l}\right).$ (37)
@@
$\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[e^{-\alpha f({x}^{i}_{0})}\right]=\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}\left[e^{-\alpha f({x}_{0})}\right]={\mathbb{E}}\left[e^{-\alpha f({x}_{0})}\right].$
@@
$<f_{*}+\frac{d\log(\alpha)}{2\alpha}+G(\alpha)$
for some function $G(\alpha)=\mathcal{O}\left(\frac{1}{\alpha}\right)$, which gives the thesis. ∎∎
Remark 3.15.
Let us consider condition (35) in the theorem above. The left hand side is larger whenever $f({x}_{0})$ is closer to $f_{*}$, while the right hand side is smaller if the initial points are closer to their average and if the error ${\cal E}_{k}$ in the function evaluation is smaller. As $\alpha$ increases, the condition becomes more restrictive, but the thesis ensures that $\operatorname*{ess\,inf}f({x}_{\infty})$ is closer to $f_{*}$.
Remark 3.16.
Let us denote with $\widehat{G}(\alpha)$ the quantity $\frac{d}{2}\frac{\log(\alpha)}{\alpha}+G(\alpha)$. We want to prove that Theorem 3.14 implies
$\mathbb{P}\left\{f(x_{\infty})-f_{*}<\widehat{G}(\alpha)\right\}>0.$
First of all, let us notice that inequality (36) is equivalent to
$\sup\big{\{}a\in\mathbb{R}\ |\ \operatorname*{ess\,inf}(f(x_{\infty})-f_{*})=a\big{\}}<\widehat{G}(\alpha)$
which in turn is equivalent to
$\sup\big{\{}a\in\mathbb{R}\ |\ \mathbb{P}\{(f(x_{\infty})-f_{*})\geq a\}=1\big{\}}<\widehat{G}(\alpha).$ (41)
Assume by contradiction that $\mathbb{P}\left\{f(x_{\infty})-f_{*}\geq\widehat{G}(\alpha)\right\}=1$. This implies
$\widehat{G}(\alpha)\in\big{\{}a\in\mathbb{R}\ |\ \mathbb{P}\{(f(x_{\infty})-f_{*})\geq a\}=1\big{\}}$
and therefore
$\sup\big{\{}a\in\mathbb{R}\ |\ \mathbb{P}\{(f(x_{\infty})-f_{*})\geq a\}=1\big{\}}\geq\widehat{G}(\alpha),$
which contradicts 41.
Let us assume the algorithm is run $R$ times, with the same parameters and the same initial distribution of the particles, and for $r=1,\dots,R$ let us denote with $S_{r}$ the random variable
$S_{r}=\begin{cases}1&\text{if }f(x_{\infty})-f_{*}<\widehat{G}(\alpha)\\ 0&\text{otherwise}\end{cases},$
and with $p$ the probability $p=\mathbb{P}\left\{S_{r}=1\right\},$ which we proved is strictly positive. We then have
$\mathbb{P}\left\{\exists\ r\in\{1,\dots,R\}\ |\ S_{r}=1\right\}=1-\mathbb{P}\left\{S_{r}=0\ \forall k=1,\dots,R\right\}=1-(1-p)^{R},$
which tends to 1 as $R$ tends to infinity. Then, increasing the number of times the algorithm is run, the probability of obtaining an accurate approximation of $f_{*}$ increases. However we notice that from Theorem 3.3 one cannot derive a lower bound for the probability $p$ of obtaining an accurate approximation of $f_{*}.$ Moreover, these inequalities concern the function values $f(x_{\infty})$ and $f_{*}$ and they do not imply in general that $x_{\infty}$ will be close to $x_{*}$.
@@ 3.5 Convergence in expectation
$+2(\gamma^{2}+\xi^{2}+(\gamma^{2}+\xi^{2})^{1/2}){\mathbb{E}}[V_{k}]^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}.$
@@
$Z_{k,1}=\frac{1}{N}\sum_{i=1}^{N}\|(\Delta x)_{k}^{i}\|^{2}=\frac{1}{N}\sum_{i=1}^{N}\left\|(\gamma I_{d}+{\bf\Xi}_{k}^{i})(x^{i}_{k}-\hat{x}^{\alpha}_{k})\right\|^{2},$ (45)
@@
$Z_{k,1}$ $\leq\frac{1}{N}\sum_{i=1}^{N}\left\|(\gamma I_{d}+{\bf\Xi}_{k}^{i})(x^{i}_{k}-{x}^{*})\right\|^{2}$ (46)
$+\frac{2}{N}\sum_{i=1}^{N}\left\|(\gamma I_{d}+{\bf\Xi}_{k}^{i})(x^{i}_{k}-{x}^{*})\right\|\left\|(\gamma I_{d}+{\bf\Xi}_{k}^{i})(\hat{x}^{\alpha}_{k}-{x}^{*})\right\|$
@@
$=(\gamma^{2}+\xi^{2}){\mathbb{E}}[\|x^{i}_{k}-{x}^{*}\|^{2}]^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}.$
In addition we note that, given $\{a_{i}\}_{i=1}^{N}$ with $a_{i}\geq 0$ for any $i$, it holds
that yields
$\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}[\|x^{i}_{k}-x^{*}\|^{2}]^{1/2}\leq\frac{1}{\sqrt{N}}\left(\sum_{i=1}^{N}{\mathbb{E}}[\|x^{i}_{k}-x^{*}\|^{2}]\right)^{1/2}={\mathbb{E}}[V_{k}]^{1/2}.$ (48)
Taking the expected value on both sides of (46) and using (48) we then have
@@
$\quad+\frac{2}{N}(\gamma^{2}+\xi^{2})\sum_{i=1}^{N}{\mathbb{E}}[\|x^{i}_{k}-{x}^{*}\|^{2}]^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}$
@@
$\leq(\gamma^{2}+\xi^{2})({\mathbb{E}}[V_{k}]+2{\mathbb{E}}[V_{k}]^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}$
@@
$+2(\gamma^{2}+\xi^{2})^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}\frac{1}{N}\sum_{i=1}^{N}{\mathbb{E}}[\|x^{i}_{k}-x^{*}\|^{2}]^{1/2}$
$\leq$ $-2\gamma{\mathbb{E}}[V_{k}]+2(\gamma^{2}+\xi^{2})^{1/2}{\mathbb{E}}[\|\hat{x}^{\alpha}_{k}-{x}^{*}\|^{2}]^{1/2}{\mathbb{E}}[V_{k}]^{1/2}.$
From (44), (49) and (51) we get the thesis. ∎∎
@@
Remark 3.18.
Assumption A6 poses condition on how well isolated the global minimizer is, in addition to the uniqueness assumption A3. The first inequality implies that in a ball around $x_{*}$, the function cannot grow too slowly. Namely, the difference $f(x)-f_{*}$ has to grow at least as $(\beta\|x-x_{*}\|)^{1/\nu}$. The second inequality ensures that, outside of a neighborhood of $x_{*}$, the function cannot take values that are too close to the minimum $f_{*}.$
The following Lemma, known in the literature as quantitative Laplace principle, provides an upper bound on $\|{x}_{k}^{\alpha}-{x}^{*}\|$, where ${x}_{k}^{\alpha}$ has been defined in (5). Consider $r\in(0,R_{0}]$ and $q>0$ such that $q+f_{r}\leq f_{\infty}$, with
@@
We distinguish the case $|{\cal I}_{k,r}|>0$ and $|{\cal I}_{k,r}|=0$. We stress that in the first case, at iteration $k$ there exists at least one particle that belongs to $B_{r}$. Part (a) of the Lemma is completely analogous to Proposition 1 in [11], while part (b) is obtained from (a) with minor changes. For completeness, a proof is provided in the appendix.
Lemma 3.19 (Quantitative Laplace Principle).
@@
    $\|{x}^{\alpha}_{k}-{x}^{*}\|\leq\frac{1}{\beta}(q+f_{r})^{\nu}+\frac{e^{-\alpha q}}{|{\cal I}_{k,r}|}\sum_{i=1}^{N}\|{x}^{i}_{k}-{x}^{*}\|,$
@@
In addition, ${\mathbb{E}}[V_{k}]\leq\mu^{k}{\mathbb{E}}[V_{0}]$ for $k=0,\dots,k_{*}.$
@@
$\|\hat{x}^{\alpha}_{\bar{k}}-{x}^{*}\|^{2}\leq$ $\|{x}^{\alpha}_{\bar{k}}-{x}^{*}\|^{2}+\|{x}^{\alpha}_{\bar{k}}-\hat{x}^{\alpha}_{\bar{k}}\|^{2}+2\|{x}^{\alpha}_{\bar{k}}-{x}^{*}\|\|{x}^{\alpha}_{\bar{k}}-\hat{x}^{\alpha}_{\bar{k}}\|$ (63)
$\leq$ $\frac{1}{4}C_{\bar{k}}^{2}+N^{2}e^{-2\alpha q}V_{\bar{k}}+Ne^{-\alpha q}V_{\bar{k}}^{1/2}C_{\bar{k}}+$
$\alpha^{2}\|{\cal E}_{\bar{k}}\|_{\infty}^{2}\sum_{s=1}^{d}\operatorname{diam}({y}^{s}_{0})^{2}\prod_{l=0}^{{\bar{k}}-1}(1-\gamma+4\max_{i=1:N}|\eta_{l,s}^{i}|)^{2}+$
$C_{\bar{k}}\alpha\|{\cal E}_{\bar{k}}\|_{\infty}\left(\sum_{s=1}^{d}\operatorname{diam}({y}^{s}_{0})^{2}\prod_{l=0}^{{\bar{k}}-1}(1-\gamma+4\max_{i=1:N}|\eta_{l,s}^{i}|)^{2}\right)^{1/2}+$
$2\alpha Ne^{-\alpha q}V_{\bar{k}}^{1/2}\|{\cal E}_{\bar{k}}\|_{\infty}\left(\sum_{s=1}^{d}\operatorname{diam}({y}^{s}_{0})^{2}\prod_{l=0}^{{\bar{k}}-1}(1-\gamma+4\max_{i=1:N}|\eta_{l,s}^{i}|)^{2}\right)^{1/2}.$
@@
which concludes the proof. ∎∎